# HateDebias: On the Diversity and Variability of Hate Speech Debiasing

### v1

@@ 2.3 Continual-learning-based Framework
To further understand the continuous debiase task and verify the benchmark dataset, we develop a continual-learning-based framework (CLF) for continuous debiasing, aimed at improving model performance and fairness by learning attribute-irrelevant features to reduce bias. Our method uses a two-stage training process: the initial training stage and the memory replay stage. Initially, the model trains on current task data to acquire new knowledge. In the memory replay stage, representative samples from previous tasks are selected to update the memory for consolidating prior knowledge. To preserve shared knowledge across tasks while integrating bias-specific knowledge, we propose a bias information regularization technique to separate bias information into a generic space and a bias-specific space, facilitating interaction and helping the model adapt to various debiasing tasks. The more detailed implementation of our framework is provided in Appendix A4.
2.3.1 Bias Information Regularization
Regularization technology separates generic and bias-specific information in hidden representations, facilitating sentence structure and context analysis. For a sentence $x$, a pre-trained language model $B(\cdot)$ produces hidden representations $h$, containing generic and bias-specific information. We then construct two networks, $G(\cdot)$ and $S(\cdot)$, to extract generic ($g$) and bias-specific ($s$) representations from hidden representations $h$. For bias-specific representation $s$, we introduce a bias-identifier prediction task that classifies examples based on their bias type using the bias-specific representation $s$. The loss function for the bias-identifier predictor is $\mathcal{L}_{task}$ (see Sec. A4.2 for detailed description).
@@ 2.3.1 Bias Information Regularization
To address catastrophic forgetting in debiasing pre-trained language models via continual learning, we propose a representation regularization approach for generic $g$ and bias-specific $s$ representations. This approach separately preserves knowledge in $g$ and $s$, promoting positive transfer and reducing forgetting. Before training on task $t$, we compute these representations for all sentences in the training set $D_{t}$ and memory buffer $M_{t}$. Using the encoder $B^{t-1}(\cdot)$ and networks $G^{t-1}(\cdot)$, $S^{t-1}(\cdot)$ from task $t-1$, we calculate $G^{t-1}(B^{t-1}(x))$ and $S^{t-1}(B^{t-1}(x))$. These are combined with task $t$’s representations for L2 norm regularization, constraining changes in both representation spaces. Thus, during the learning in debiasing task $t$, we impose two regularization losses respectively:
$\mathcal{L}_{g}=E_{x\in D_{t}\cup M_{t}}||G^{t-1}(B^{t-1}(x))-G^{t}(B^{t}(x))||_{2},$ (1)
$\mathcal{L}_{s}=E_{x\in D_{t}\cup M_{t}}||S^{t-1}(B^{t-1}(x))-S^{t}(B^{t}(x))||_{2},$ (2)
@@
2.3.2 Memory Replay
To efficiently use the memory buffer $M$ while maintaining balanced replay, we selectively store a small number of samples from the previous debiasing task’s training set. We explore strategies to recall specific samples to mitigate model bias, which arises from performance variations across groups defined by a sensitive attribute. Replaying low-accuracy samples, particularly those associated with a sensitive attribute, can effectively reduce bias. In the $t$-th debiasing stage, before fine-tuning and debiasing the model, we select the corpus $D_{r}$ for stage $r(r\textless t)$. First, the class predictor $f_{cls}(\cdot)$ is used to predict $D_{r}$, and the probability list $P$ corresponding to the true label of each sample is obtained. We set the number of retained samples as $\gamma|D_{r}|$, and obtain the corresponding probability threshold as $\beta$, then the samples retained in stage $r$ are:
@@ 2.3.2 Memory Replay
2.3.3 Overall Objective
When performing the $t$-th debiasing task, we simultaneously leverage the generic representation $g$ and the bias-specific representation $s$ for text classification. To optimize model performance, we minimize the cross-entropy loss $\mathcal{L}_{cls}$ (see Sec. A4.3 for detailed description). In addition, for a given sentence $x$, we adopt the given debiaser $H_{t}(\cdot)$ to mitigate attribute $A_{t}$ in the text. The debiaser $H_{t}(\cdot)$ is designed to be highly flexible and extensible, allowing for modifications as needed. In our experiments, we utilize four debiasing methods as the backbones of the debiaser $H_{t}(\cdot)$, namely FGM [15], PGD [16], ATC [17, 18, 19] and CL [20, 21, 22]. We define the debiaser’s loss as $\mathcal{L}_{d}$ (see Sec. A4.4 for detailed description). The final loss of our continual debiasing model is:
@@ 2.3.3 Overall Objective
3 Experiment
We design multiple sets of experiments to validate the effectiveness of our proposed model on the proposed dataset. First, we evaluate that existing methods can hardly achieve the ideal performance in the multiple biases while the proposed method can solve this problem, which is shown in 3.2 and Sec. 3.3. Sequentially, to further evaluate that our method can achieve stable performance under the continuous scenarios of our datasets, we further consider different numbers of biases in the continuous scenarios, which is shown in Sec. 3.4. Moreover, we evaluate that our method can avoid catastrophic forgetting as shown in Sec. 3.5. Besides, we conduct ablation experiments, parameter search experiments, statistical significance tests, explore the effects of different sequence orders, and different pre-trained language models. Please refer to more experiment results in Appendix A7.
3.1 Setups
@@
We consider seven baselines: Fine-tune [23], Experience Replay (ER) [24, 11], IDBR [25], Multi-task Learning (MTL) [26], Two-level Knowledge Distillation (TKD) [27], Representation Projection Invariance (REPINA) [28] and (Instance-wise Relation Distillation (IRD) [29] to combine debiasing methods to verify the effectiveness of various strategies and our proposed method. Given that our experimental setting involves attribute-incremental continual learning, we transform the attribute-incremental setting into a debiasing-task-incremental setting.
@@
Table 4: The optimal hyperparameter values of each debiasing method.
@@
All experiments are carried out based on PyTorch and an A30 with 24 GB of memory. We design our framework based on Transformers. Furthermore, in terms of the pre-trained language model, we choose the bert-base-uncased model [30] as the backbone of our framework. We set the same hyper-parameters with a fixed initialization seed for our model’s training, where the batch size is 32 and the feature dimension is 768. The max length of the input samples is 32. The BERT encoder with a learning rate of 5e-5 is optimized utilizing Adam. For the MTL model, we set the training epoch to 10, while for other continual learning methods, we set the number of epochs to 5 for each stage. We perform a grid search on the hyperparameters in the model. The optimal hyperparameter values of each debiasing method are shown in Table 4. In order to avoid the contingency caused by the order of debiased attributes, we conduct experiments with $A_{4}^{4}=24$ combinations of 4 attributes.
@@
Formally speaking, in the continuous debiasing task, we require that a certain type of bias cannot be detected in the model after the debiasing function $MBais(\cdot)$. Unlike traditional task-incremental continual learning tasks, we perform attribute-incremental continual learning on the same downstream task. We adopt the accuracy (Acc) and macro-average $F_{1}$ (F1-macro) value to evaluate the performance of models in the downstream task. To evaluate group fairness, we measure the equality differences of false positive rates (FPR) for each bias. In terms of the overall fairness of the model, the average false positive rate of four biases is calculated to measure, which is defined as the averaged attribute bias (AAB). Since there are 24 sequence settings in the attribute-incremental continual learning task, for continual-learning-based methods, we calculate the average of these metrics for all task sequences. For each metric, we also calculate the variance.
Table 5: Main results. The columns and rows are the evaluation metrics and the names of each method respectively. Values in brackets are the variances of the metrics. For the static debiasing method, since there are no multiple sequence experiments, we denote the variance as 0.00. MTL represents a static debiasing approach that addresses multiple sensitive attributes concurrently. Fine-tune constitutes the most straightforward experimental setup for continual debiasing tasks. Furthermore, ER, IDBR, REPINA, IRD and TKD are prevalent methods in continuous learning for continual debiasing tasks.
Method Debiasing Method Performance Fairness
Acc $\uparrow$ F1-macro $\uparrow$ FPR(Gender) $\downarrow$ FPR(Age) $\downarrow$ FPR(Country) $\downarrow$ FPR(Ethnicity) $\downarrow$ AAB $\downarrow$
MTL [26] No-debias 85.90(0.00) 83.00(0.00) 4.77(0.00) 7.77(0.00) 5.35(0.00) 4.43(0.00) 5.58(0.00)
PGD 86.79(0.00) 83.59(0.00) 3.05(0.00) 7.37(0.00) 3.21(0.00) 3.87(0.00) 4.37(0.00)
FGM 87.19(0.00) 83.99(0.00) 3.92(0.00) 6.05(0.00) 4.97(0.00) 4.18(0.00) 4.78(0.00)
ATC 85.48(0.00) 82.32(0.00) 6.11(0.00) 8.28(0.00) 5.55(0.00) 3.87(0.00) 5.95(0.00)
CL 88.12(0.00) 84.38(0.00) 4.76(0.00) 7.04(0.00) 6.26(0.00) 5.17(0.00) 5.81(0.00)
Fine-tune [23] No-debias 86.73(2.92) 83.51(3.23) 5.55(0.91) 8.67(2.16) 5.97(1.67) 4.51(1.10) 6.17(0.65)
PGD 87.14(3.11) 84.23(3.42) 5.31(0.68) 8.67(4.52) 5.56(0.68) 4.06(0.60) 5.90(0.75)
FGM 87.07(3.31) 84.16(3.68) 5.43(1.22) 8.86(5.80) 5.74(0.93) 4.39(0.52) 6.10(1.01)
ATC 88.14(1.34) 85.17(1.36) 5.95(0.60) 10.02(2.00) 6.56(1.19) 4.85(0.88) 6.84(0.30)
CL 86.59(2.72) 83.41(3.32) 5.15(0.89) 8.85(1.70) 6.19(0.87) 4.28(0.83) 6.12(0.37)
ER [11] PGD 88.68(0.32) 85.64(0.37) 5.38(0.70) 8.63(0.70) 5.43(0.86) 4.64(0.82) 6.02(0.31)
FGM 88.64(0.55) 85.60(0.68) 5.53(0.23) 9.04(0.74) 5.27(0.42) 4.83(0.87) 6.17(0.21)
ATC 87.17(1.72) 83.83(1.61) 5.10(0.62) 8.58(2.41) 6.32(1.52) 4.79(1.13) 6.20(0.37)
CL 87.19(0.73) 84.05(0.88) 4.96(0.93) 8.99(2.43) 6.25(0.82) 4.83(1.59) 6.26(0.47)
IDBR [25] ATC 88.74(0.87) 85.81(1.40) 5.30(1.34) 9.45(1.99) 6.53(1.30) 4.76(0.48) 6.51(0.51)
CL 87.57(0.80) 84.40(0.90) 4.33(1.18) 8.18(1.43) 5.32(1.24) 4.70(1.21) 5.63(0.64)
REPINA [28] ATC 87.42(0.97) 84.27(1.21) 5.71(0.79) 9.72(3.06) 5.83(0.45) 4.56(1.00) 6.46(0.37)
CL 88.50(0.84) 85.29(1.31) 6.05(1.52) 9.73(4.36) 6.64(1.40) 4.40(1.21) 6.70(1.12)
IRD [29] ATC 88.70(0.77) 85.79(1.00) 5.66(0.70) 9.94(1.37) 6.74(1.10) 5.28(0.62) 6.90(0.45)
CL 87.11(0.97) 83.84(1.19) 4.41(2.18) 8.02(3.92) 5.40(1.37) 4.49(1.65) 5.58(1.14)
TKD [27] ATC 88.86(1.07) 85.96(1.34) 5.58(0.79) 9.66(0.87) 6.37(1.23) 4.81(1.07) 6.60(0.26)
CL 87.02(1.51) 83.70(1.61) 4.58(1.86) 8.42(2.00) 5.33(0.05) 4.51(1.30) 5.71(0.91)
CLF PGD 87.16(2.02) 84.01(2.16) 5.79(0.93) 8.64(1.48) 4.96(0.56) 4.38(0.89) 5.94(0.46)
FGM 86.37(7.63) 83.35(7.12) 5.78(0.93) 8.00(4.16) 4.57(0.52) 4.35(2.65) 5.67(0.70)
ATC 83.85(13.81) 80.51(13.78) 3.90(2.51) 7.33(5.08) 4.06(2.82) 4.22(1.26) 4.88(1.28)
CL 83.94(6.91) 80.31(6.42) 2.79(2.82) 6.21(2.60) 3.61(2.65) 4.84(1.60) 4.36(1.45)
3.2 Debiasing Performance Forgetting of Continuous Debiasing

### v2

@@ IV-C Continual-learning-based Framework
To further understand the continuous debiase task and verify the benchmark dataset, we develop a continual-learning-based framework (CLF) for continuous debiasing, aimed at improving model performance and fairness by learning attribute-irrelevant features to reduce bias. Our method uses a two-stage training process: the initial training stage and the memory replay stage. Initially, the model trains on current task data to acquire new knowledge. In the memory replay stage, representative samples from previous tasks are selected to update the memory for consolidating prior knowledge. To preserve shared knowledge across tasks while integrating bias-specific knowledge, we propose a bias information regularization technique to separate bias information into a generic space and a bias-specific space, facilitating interaction and helping the model adapt to various debiasing tasks. The pseudocode of the implementation process of our proposed continual-learning-based debiasing framework is shown in Algorithm 1.
Algorithm 1 CLF
2:Trained language model encoder $LM^{n}$
3:Load pre-trained language model $LM^{0}$
4:for i=1,,ni=1,...,n do
7:  if $i\geq 2$ then
8:    Calculate $\mathcal{L}_{g}$ and $\mathcal{L}_{s}$ with $LM^{i}$ (Eq. 2-3) $\triangleright$ Bias Information Regularization
10:    $M_{t}$ = []
11:    for r=1,,ir=1,...,i do $\triangleright$ Memory Replay
12:     $M_{t}\leftarrow\gamma|D_{r}|$ (Eq. 6)
13:     $M_{t}\leftarrow M_{t}\cup M_{t}^{r}$
14:    end for
15:  end if
17:  Update model parameters with $\mathcal{L}$
18:end for
IV-C1 Overview
We present a continual-learning-based framework for continual debiasing tasks, aiming to enhance the model’s performance and fairness by promoting the learning of attribute-irrelevant features to effectively mitigate bias. For a given task sequence $\mathcal{T}_{\textit{debias}}=\{\operatorname{T_{1}},\cdots,\operatorname{T_{n}}\}$, our proposed method is adopted to the two-stage training backbone, in terms of the initial training stage and the memory replay stage, where the replay-based method can effectively alleviate the classifier bias. During the initial training stage, the model performs training solely on the current task data, allowing the learning of new task-related knowledge. Subsequently, in the memory replay stage, a subset consisting of representative samples from the previous task data is selected to update the memory. The model is further retrained using the balanced memory bank, facilitating the retrieval and consolidation of previously acquired knowledge. Specifically, after training on (t1)(t-1)-th $(t\geq 2)$ task $\operatorname{T}_{t-1}$, $\gamma|D_{t-1}|$ examples are sampled from the training set $D_{t-1}$ into the memory buffer $M$, where $0\leq\gamma\leq 1$ is the store ratio. Data from $M$ is then merged with the next training set $D_{t}$ when conducting the learning of task $t$. What’s more, in the process of continual learning, an important aspect arises from the presence of shared knowledge across different debiasing tasks. Hence, the model must constantly adapt to the specific knowledge related to each debiasing task while avoiding forgetting the old knowledge throughout the learning process. To tackle this challenge, we disentangle sensitive attribute information from raw text representation to obtain a generic space and bias-specific space, introducing an bias information regularization technique to allow interaction between two spaces. The regularization approach aims to preserve shared knowledge while enabling the model to effectively incorporate bias-specific knowledge for different debiasing tasks.
IV-C2 Bias Information Regularization
The operation of disentangled enables the separation of generic and bias-specific information in the hidden representations, facilitating a more refined analysis and understanding of the sentence’s underlying structure and context. The disentangled concept we define allows the model to learn generic information about downstream tasks and specific information about sensitive attributes, respectively. For a given sentence $x$, a pre-trained language model $LM(\cdot)$ is employed to derive the hidden representations $h$, encompassing both generic and bias-specific information. Subsequently, we design two disentangled networks, represented by $G(\cdot)$ and $S(\cdot)$, which play a vital role in extracting the generic representation $g$ and sensitive bias-specific representation $s$ from the previously obtained hidden representations $h$. $G(\cdot)$ and $S(\cdot)$ refer to two distinct feed-forward networks. The dimensions of the representations produced by $G(\cdot)$ and $S(\cdot)$ are identical to the hidden representations $h$. This systematic process enables the generic and bias-specific information to disentangle, facilitating a more refined analysis and understanding of the sentence’s underlying structure and context. For the dynamic debiasing task of sensitive attributes, we train a classifier by exploiting information from two representations and allowing different representation spaces to be changed to different extents, maintaining the ability of the models when debiasing the sensitive attributes.
Generic Representation
Bias-specific Representation
$\mathcal{L}_{task}=E_{(x,z)\in D_{t}\cup M_{t}}\mathcal{L}_{CE}(f_{task}(S(LM(x)),z),$ (1)
where $z$ is the corresponding debiasing task id for $x$, $L_{CE}(\cdot)$ is the cross entropy loss function and $LM(\cdot)$ is the pre-trained model for text encoding in our framework. The memory buffer, denoted as $M_{t}$, retains samples from the training sets of prior tasks. This auxiliary setup serves as a valuable mechanism for enhancing task-specific learning and contributes to the robustness and adaptability of the model across different tasks.
Representation Regularization
@@ IV-C2 Bias Information Regularization
To cope with catastrophic forget during continually debiasing pre-trained language models via continual learning, we propose a regularization approach applied to both generic representations $g$ and bias-specific representations $s$, which preserves knowledge in distinct spaces separately to encourage positive transfer and mitigate forgetting.
We design two disentangled networks $G^{t-1}(\cdot)$ and $S^{t-1}(\cdot)$. We compute the generic representation $g$ via $G^{t-1}(LM^{t-1}(x))$ and the bias-specific representation $s$ via $S^{t-1}(LM^{t-1}(x))$ based on $D_{t-1}$, thereby preserving generic knowledge from the prior models to alleviate forgetting. Then, for the debiasing task $t$, we obtain the generic representations and bias-specific representations for all sentences within the training set $D_{t}$ and the memory buffer $M_{t}$ ($M_{t}\subseteq M$). Finally, the general and bias-specific representations from task t-1 are retained together with the representations obtained in task $t$ for L2 norm regularization, thereby limiting the extent of updates to the general representation space and bias-specific representation space. Thus, during the learning for training pairs in the debiasing task $t$, we impose two regularization losses respectively:
[…]
@@
IV-C3 Text Classification
where $y$ denotes the corresponding class label assigned to input $x$, and $f_{cls}(\cdot)$ represents the class predictor function. The symbol $\oplus$” signifies the concatenation operation applied to the two representations under consideration.
IV-C4 Text Debiasing
In the $t$-th debiasing stage, the model is aimed at debiasing the sensitive attribute $A_{t}$. Specifically, for a given sentence $x$, we adopt the given debiaser $H_{t}(\cdot)$ to mitigate sensitive attribute $A_{t}$ in the text. The debiaser $H_{t}(\cdot)$ is designed to be highly flexible and extensible, allowing for modifications as needed. In our experiments, we utilize four debiasing methods as the backbones of the debiaser $H_{t}(\cdot)$, namely FGM [40], PGD [41], ATC [15] and CL [42]. More details are presented in Section V-A. We minimize the debiaser’s loss:
$\mathcal{L}_{d}=E_{(x,y)\in D_{t}\cup M_{t}}\mathcal{L}_{d}(H_{t}(x)).$ (5)
IV-C5 Memory Replay
@@
IV-C6 Overall Objective
The final loss of our dynamic debiasing model is:
@@ IV-C6 Overall Objective
IV-C7 Generalizability of Our Framework
V Experiment
We design multiple sets of experiments to validate the effectiveness of our proposed model on the proposed dataset. First, we evaluate that existing methods can hardly achieve the ideal performance in the multiple biases while the proposed method can solve this problem, which is shown in V-B and Sec. V-C. Sequentially, to further evaluate that our method can achieve stable performance under the continuous scenarios of our datasets, we further consider different numbers of biases in the continuous scenarios, which is shown in Sec. V-D. Moreover, we evaluate that our method can avoid catastrophic forgetting as shown in Sec. V-E. Besides, we conduct ablation experiments, parameter search experiments, statistical significance tests, explore the effects of different sequence orders, and different pre-trained language models.
V-A Setups
Debiasing Methods
We conduct experiments based on four debiasing methods in the continuous debiasing task. Fast Gradient Method (FGM) [40] follows the adversarial training method to add perturbation to debiasing model representation module. Projected Gradient Descent (PGD) [41] is another unsupervised adversarial learning method, which adds an iterative attack that differs from FGM. Adversarial Training Classifiers (ATC) tend to perform supervised adversarial training [15, 43, 44] to predict sensitive attributes for adversarial learning, thereby blurring the model’s identification of sensitive attributes. Contrastive Learning (CL) [42] aims at pulling identical entailment pairs along opposite sensitive attribute directions closer, which serves as an intermediate pre-training approach to mitigate biases in contextual representations.
@@ V-A Setups
We compare with the following baselines to verify the effectiveness of our proposed method. Fine-tune [45] modifies the parameters of the pre-trained language model sequentially. Experience Replay (ER) [11, 26] is a memory-based technique that involves retaining a limited subset of samples from prior tasks and subsequently recalling them during training. IDBR [46] incorporates information disentanglement regularization into the encoding process. Multi-task Learning (MTL) [47] trains the model on all tasks simultaneously to avoid the problem of catastrophic forgetting, as a static debiasing method. Two-level Knowledge Distillation (TKD) [48] is a technique designed to improve the continual text classification. It involves transferring knowledge at the feature level and the output level. Representation Projection Invariance (REPINA) [49] aims to preserve the information content of representations by discouraging undesirable changes. Instance-wise Relation Distillation (IRD) [50] is a regularization technique used in continual learning to mitigate the forgetting of previously learned knowledge when the model learns new tasks. Since our experimental setting involves attribute-incremental continual learning, we transform the attribute-incremental setting into a debiasing-task-incremental setting to enable the task-incremental continual learning method to work on our dataset.
@@
TABLE IV: The optimal hyperparameter values of each debiasing method.
@@
TABLE V: Main results. The columns and rows are the evaluation metrics and the names of each method respectively. Values in brackets are the variances of the metrics. For the static debiasing method, since there are no multiple sequence experiments, we denote the variance as 0.00. MTL represents a static debiasing approach that addresses multiple sensitive attributes concurrently. Fine-tune constitutes the most straightforward experimental setup for continual debiasing tasks. Furthermore, ER, IDBR, REPINA, IRD and TKD are prevalent methods in continuous learning for continual debiasing tasks.
Method Debiasing Method Performance Fairness DTO
Acc $\uparrow$ F1-macro $\uparrow$ FPED(Gender) $\downarrow$ FPED(Age) $\downarrow$ FPED(Country) $\downarrow$ FPED(Ethnicity) $\downarrow$ AAB $\downarrow$
MTL [47]
No-debias 85.90(0.00) 83.00(0.00) 4.77(0.00) 7.77(0.00) 5.35(0.00) 4.43(0.00) 5.58(0.00) 0.28
PGD 86.79(0.00) 83.59(0.00) 3.05(0.00) 7.37(0.00) 3.21(0.00) 3.87(0.00) 4.37(0.00) 0.02
FGM 87.19(0.00) 83.99(0.00) 3.92(0.00) 6.05(0.00) 4.97(0.00) 4.18(0.00) 4.78(0.00) 0.10
ATC 85.48(0.00) 82.32(0.00) 6.11(0.00) 8.28(0.00) 5.55(0.00) 3.87(0.00) 5.95(0.00) 0.37
CL 88.12(0.00) 84.38(0.00) 4.76(0.00) 7.04(0.00) 6.26(0.00) 5.17(0.00) 5.81(0.00) 0.33
Fine-tune [45]
No-debias 86.73(2.92) 83.51(3.23) 5.55(0.91) 8.67(2.16) 5.97(1.67) 4.51(1.10) 6.17(0.65) 0.42
PGD 87.14(3.11) 84.23(3.42) 5.31(0.68) 8.67(4.52) 5.56(0.68) 4.06(0.60) 5.90(0.75) 0.35
FGM 87.07(3.31) 84.16(3.68) 5.43(1.22) 8.86(5.80) 5.74(0.93) 4.39(0.52) 6.10(1.01) 0.40
ATC 88.14(1.34) 85.17(1.36) 5.95(0.60) 10.02(2.00) 6.56(1.19) 4.85(0.88) 6.84(0.30) 0.57
CL 86.59(2.72) 83.41(3.32) 5.15(0.89) 8.85(1.70) 6.19(0.87) 4.28(0.83) 6.12(0.37) 0.40
ER [11]
PGD 88.68(0.32) 85.64(0.37) 5.38(0.70) 8.63(0.70) 5.43(0.86) 4.64(0.82) 6.02(0.31) 0.38
FGM 88.64(0.55) 85.60(0.68) 5.53(0.23) 9.04(0.74) 5.27(0.42) 4.83(0.87) 6.17(0.21) 0.42
ATC 87.17(1.72) 83.83(1.61) 5.10(0.62) 8.58(2.41) 6.32(1.52) 4.79(1.13) 6.20(0.37) 0.42
CL 87.19(0.73) 84.05(0.88) 4.96(0.93) 8.99(2.43) 6.25(0.82) 4.83(1.59) 6.26(0.47) 0.44
IDBR [46]
ATC 88.74(0.87) 85.81(1.40) 5.30(1.34) 9.45(1.99) 6.53(1.30) 4.76(0.48) 6.51(0.51) 0.49
CL 87.57(0.80) 84.40(0.90) 4.33(1.18) 8.18(1.43) 5.32(1.24) 4.70(1.21) 5.63(0.64) 0.29
REPINA [49]
ATC 87.42(0.97) 84.27(1.21) 5.71(0.79) 9.72(3.06) 5.83(0.45) 4.56(1.00) 6.46(0.37) 0.48
CL 88.50(0.84) 85.29(1.31) 6.05(1.52) 9.73(4.36) 6.64(1.40) 4.40(1.21) 6.70(1.12) 0.54
IRD [50]
ATC 88.70(0.77) 85.79(1.00) 5.66(0.70) 9.94(1.37) 6.74(1.10) 5.28(0.62) 6.90(0.45) 0.58
CL 87.11(0.97) 83.84(1.19) 4.41(2.18) 8.02(3.92) 5.40(1.37) 4.49(1.65) 5.58(1.14) 0.28
TKD [48]
ATC 88.86(1.07) 85.96(1.34) 5.58(0.79) 9.66(0.87) 6.37(1.23) 4.81(1.07) 6.60(0.26) 0.51
CL 87.02(1.51) 83.70(1.61) 4.58(1.86) 8.42(2.00) 5.33(0.05) 4.51(1.30) 5.71(0.91) 0.31
CLF PGD 87.16(2.02) 84.01(2.16) 5.79(0.93) 8.64(1.48) 4.96(0.56) 4.38(0.89) 5.94(0.46) 0.36
FGM 86.37(7.63) 83.35(7.12) 5.78(0.93) 8.00(4.16) 4.57(0.52) 4.35(2.65) 5.67(0.70) 0.30
ATC 83.85(13.81) 80.51(13.78) 3.90(2.51) 7.33(5.08) 4.06(2.82) 4.22(1.26) 4.88(1.28) 0.12
CL 83.94(6.91) 80.31(6.42) 2.79(2.82) 6.21(2.60) 3.61(2.65) 4.84(1.60) 4.36(1.45) 0.06
We conduct all experiments based on PyTorch and an A30 with 24 GB of memory. We design our framework based on Transformers. Furthermore, we choose the bert-base-uncased model [51] as the backbone of our framework. We set the same hyper-parameters with a fixed initialization seed for our model’s training, where the batch size is 32 and the feature dimension is 768. The max length of the input samples is 32. The BERT encoder with a learning rate of 5e-5 is optimized utilizing Adam. For the MTL model, we set the training epoch to 10, while for other continual learning methods, we set the number of epochs to 5 for each stage. We perform a grid search on the hyperparameters in the model. The optimal hyperparameter values of each debiasing method are shown in Table IV. In order to avoid the contingency caused by the order of debiased attributes, we conduct experiments with $A_{4}^{4}=24$ combinations of 4 attributes.
@@
Unlike traditional task-incremental continual learning tasks, we perform attribute-incremental continual learning on the hate speech detection task to debias for a specific attribute. We adopt the accuracy (Acc) and macro-average $F_{1}$ (F1-macro) value to evaluate the performance of models on the hate speech detection task. To evaluate model fairness, we measure the equality differences of false positive rates (FPED) across different groups for a specific attribute. As formulated in Eq. 8, FPED involves aggregating the differences between the false positive rates of each subgroup $FPR_{d}$ and the overall false positive rate $FPR$ within a demographic factor $D$. Since there are 24 sequence settings for each attribute in the attribute-incremental continual learning task, we calculate the average FPED across all task sequences for each attribute in our continual-learning-based framework. In terms of the overall fairness of the model, we compute the average FPED of four attribute biases to measure the average attribute bias (AAB). Since there are 24 sequence settings in the attribute-incremental continual learning task, we calculate the AAB for all task sequences for our continual-learning-based framework. For each metric, we also calculate the variance. Moreover, we further introduce the DTO metric [52, 53] to measure the trade-off between the fairness improvements and performance impacts. A lower DTO value indicates a more effective trade-off.
$FPED=\sum_{d\in D}\left|FPR_{d}-FPR\right|.$ (8)
V-B Debiasing Performance Forgetting of Continuous Debias- ing